From a pair of simultaneous equations for each situation below, and use them to solve the problem. Let the unknown numbers be x and y.
Q. The combined ages of two people is 34. If one person is 6 years younger than other, find their ages.

Answer by stanbon(75887)

Form a pair of simultaneous equations for each situation below, and use them to solve the problem. Let the unknown numbers be x and y.
---------
Q. The combined ages of two people is 34. If one person is 6 years younger than other, find their ages.
----------
x + y = 34
x = y-6
------
Substitute for "x" and solve for "y":
y-6 + y = 34
2y = 40
y = 20 yrs
-----
x = y-6 = 14 yrs
====
Cheers,
Stan H.
-------------